OpenStudy (anonymous):

solve 3x^2+18x+15=0 by completing the square

OpenStudy (anonymous):

@surjithayer

OpenStudy (anonymous):

@AccessDenied

OpenStudy (anonymous):

hm

OpenStudy (jdoe0001):

do you know what a perfect square trinomial is?

OpenStudy (anonymous):

1.take common 3 2.add the square of half the co-efficient of x.

OpenStudy (anonymous):

im confused

OpenStudy (anonymous):

\[3\left( x^2+6x \right)+5=0,3\left( x^2+6x+9-9 \right)+5=0\] \[3\left( x^2+6x+9 \right)-27+5=0\] \[3\left( x+3 \right)^2-22=0\] \[\left( x+3 \right)^2=\frac{ 22 }{ 3 },x+3=\pm \sqrt{\frac{ 22 }{ 3 }}\] find x

OpenStudy (anonymous):

im still confused

OpenStudy (mathmale):

solve 3x^2+18x+15=0 by completing the square: Let's go thru this quickly, with your involvement. Later, we could discuss the why's. 3 is a factor common to all four terms in 3x^2+18x+15=0. Please factor out the 3 right now: 3( ??? )

OpenStudy (mathmale):

15/3 = ?

OpenStudy (anonymous):

15 x 3 or 15 divided by 3?

OpenStudy (mathmale):

I was just trying to help you get started on the necessary factoring.\[\frac{ 15 }{ 3 }=?\]

OpenStudy (anonymous):

5

OpenStudy (mathmale):

Right. and (3x^2)/3 = ?

OpenStudy (anonymous):

3?

OpenStudy (mathmale):

\[\frac{ 3x^2 }{ 3 }=?\]Hint: just cancel the 3s.

OpenStudy (anonymous):

x^2

OpenStudy (mathmale):

Right. and \[\frac{ 18x }{ 3 }=?\]

OpenStudy (anonymous):

6

OpenStudy (mathmale):

Actually, that'd be 6x. Then 3x^2+18x+15=0 has been factored: 3(x^2 + 6x +5)=0. Look carefully. Does this make sense to you?

OpenStudy (anonymous):

yes

OpenStudy (mathmale):

Great. Now, we're going to ignore that coefficient 3 for a while, and focus only on x^2 + 6x +5. We're going to do something to x^2 + 6x + 5 called "completing the square." I won't explain a lot right now, but rather will just ask you to do this or that; we'll go into the reasons later if you're interested.

OpenStudy (mathmale):

Re-write x^2 + 6x + 5 as x^2 + 6x + 5 Now take the coefficient of x (which is 6); halve it. What do you get? Half of 6 is what?

OpenStudy (anonymous):

3

OpenStudy (mathmale):

Good. Now please square that 3. 3^2 = ?

OpenStudy (anonymous):

9

OpenStudy (mathmale):

Right you are! Now take another look at x^2 + 6x + 5 Add your 9 to the right of the 6x term; then subtract 9 from the 1st 9.

OpenStudy (anonymous):

huh?

OpenStudy (mathmale):